OpenStudy (anonymous):

Solve the integro-differential for y(t) y(t) + 3*int_{0}^{t} y(v)*sin(t-v)dv = t in case the integral thinger doesn't work y(t_ +3* the integral of [y(v)*sin(t-v)dv] = t lower limit is 0. upper limit is t

OpenStudy (anonymous):

fail. equations hurr: \[y(t) + 3*\int\limits_{0}^{t} y(v)\sin(t-v)dv = t\]

OpenStudy (anonymous):

subject is differential equations, section is convolutions

OpenStudy (anonymous):

god, not that.

OpenStudy (anonymous):

you supposed to be using laplace transform?

OpenStudy (anonymous):

yessir

OpenStudy (anonymous):

have a test in 5 hours lol. last thing i need to understand before i feel ready.

OpenStudy (anonymous):

I think it's just \[Y(s) +\frac{ 3Y(s) }{ s^2+1} = \frac{ 1 }{ s^2 }\]

OpenStudy (anonymous):

i think there's supposed to be a sin(something) + something

OpenStudy (anonymous):

I'm moving on in about 3 minutes lol. still get to do power and taylor series and dirac delta functions. yayyyyy lol

OpenStudy (anonymous):

solve for Y(s) and inverse transform...

OpenStudy (anonymous):

sec

OpenStudy (anonymous):

gahhhh. im at y(t) = t - sin(t)*y(t) the * is the convolution symbol

OpenStudy (anonymous):

hmm I got (1/4)sin(2t) + t*(1/4)sin(2t)) the * is the convolution symbol

OpenStudy (anonymous):

been a while since I've done these, though...

OpenStudy (anonymous):

so \[Y = \frac{ 1 }{ s^2 } - (\frac{ 1 }{ s^2+1 })*(Y)\] if that makes any sense

OpenStudy (anonymous):

i have a feeling this test is gonna tank my grade haha

OpenStudy (anonymous):

\[Y(s) = \frac{ s^2+1 }{ s^2(s^2+4) }\] did you get that?

OpenStudy (anonymous):

nada

OpenStudy (anonymous):

i dont know what im getting yet lol. lemme roll w/ that for a sec tho

OpenStudy (anonymous):

so using that i got 3sin(2t)/8 + t/4

OpenStudy (anonymous):

how did you get to that point?

OpenStudy (anonymous):

solved Y(s)+3Y(s)/(s^2+1)=1/s^2 for Y(s)

OpenStudy (anonymous):

but how did you get 3sin(2t)/8 + t/4 ?

OpenStudy (anonymous):

partial fraction decomp and then inverse LT

OpenStudy (anonymous):

how did you get to this point lol solved Y(s)+3Y(s)/(s^2+1)=1/s^2 for Y(s)

OpenStudy (anonymous):

convolution in time domain is multiplication in s domain so L(sin(t)) *L(y(t))

OpenStudy (anonymous):

yeah I see... would've have thought that wouldn't * work here, so I didn't try it..

OpenStudy (anonymous):

kkkkkkk got it :) <3

OpenStudy (anonymous):

thanks for the help, onto power and taylor series!!! lol